OpenStudy (anonymous):

Can you check my answer for me please: Essay, show all work. Consider that the age, x, of a unicorn in human equivalent years can be given by the formula f(x)= -0001518x^4 + 0.067326x^3 - 1.4367x^2 +12.46x +2.914. When a unicorn is 2.5 years old, what is its age in human equivalent years? What about when it is 12 years old? When a unicorn is 2.5 years old, the age in human equivalent years is: f(x)= -0.001518(2.5)^4 +0.067326(2.5)^3 - 1.4367(2.5)^2 +12.46(2.5) +2.914 f(x)= -0.003795^4 + 0.168315^3 - 3.59175^2 +31.15 +2.914 f(x)= -0.003795^4 + 0.168315^3 - 3.59175^2 +34.064 When a unicorn i

OpenStudy (anonymous):

When a unicorn is 12 years old, the age in human equivalent years is: f(x)= -0.001518(12)^4 +0.067326(12)^3 - 1.4367(12)^2 +12.46(12) +2.914 f(x)= -0.018216^4 +0.807912^3 -17.2404^2 +149.52 +2.914 f(x)= -0.018216^4 +0.807912^3 -17.2404^2 +152.434

OpenStudy (anonymous):

Well, I guess no one wants to answer this one?

OpenStudy (anonymous):

the power on the x variable does not transfer over to the decimal point, therefore: -0.001518(12)^4=-0.001518(20736)=-37.477248 and so on for all your powers. use this property to recheck your answers. also try a calculator, those calculations are a pain by hand